need a bode plot for gain and phase
(145) (5) (49) | - (9)
Answer #1

Given transfer function is G(s)= \frac{1}{(s+1)(s+1)(s+10)}=\frac{1}{(s+1)^{2}(s+10)}

System has 1 input and one out put.

ordr of system is 3

number of zeros =0

number of poles are 3 that are s=-1 (order 2 pole) s= -10 (simple pole)